AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 374     Entry time: Thu Aug 11 09:25:13 2022
Author: JM, TD 
Subject: Noise tests 


Moved pneumatic motor in and out several revolutions

Biased detector up to 100 V, I ~ 12.4 uA ambient temp c. 21 deg C

Moving detector in increased leakge current to ~14.5 uA, moving detector out dropped leakage current back to ~12.4 uA

aida05 - pulser width 21 chn (slightly lower than yesterday)

aida01 - pulser width 124 chn

aida02 - double peak (is not currently grounded)

aida03 - 183 chn

aida04 - 212 chn

 

Attachment 1     1.8.L spectra p+n FEE64s

Attachment 2      per FEE64 rate spectra

Attachment 3      aida05 1.8.L, 1.8.W and rate spectra

Attachment 4      FEE64 temperatures - OK

Attachment 5      per FEE64 ADC data event stats - rates for aida01-aida04 significabtly lower than yesterday presumably due to detector bias -100V

Attachments 6 & 7 per FEE64 1.8.W spectra

Attachment 8      1.8.L spectra for n+n FEE64s

regrounded aida02 back to cryring ground - required removing star configuration of all fee grounds and making a new

aida05 - pulser width 30 chn 

aida01 - pulser width 300 chn

aida02 - 220 -> double peak gone 

aida03- 190

aida04 - 200

Perhaps star configuration isn't as tight as before - re tighten star section

Re tightened section -> no impact on pulser widths

Moved detector in using pneumatic -> no impact on pulser widths

Added a grounding cable from the 19 inch rack to the same CRYRING ground that the fees are connected too. Stats around 2-3k, slow comparator = 0x64, bias=100v, I=12.18 uA

aida01  ~ 260

aida02 ~ 177

aida03 ~ 178

aida04 ~ 187

aida05 ~28

Removed CRYRING ground. Fees and rack connected together still. 

no change

Bias supply module is floating

ion pumps off, gauge off -> no change in waveforms, leakage current dropped to 8.7 uA

aida01 ~ 255

aida02 ~ 176

aida05 ~ 26

Removed all vacuum systems and motors power from 19 inch rack power. Connected to the rack are now: NIM bin, network switch, 2 inputs to mains relay, raspberrypi, pi dlink usb connector, VME crack

aida01 ~ 220

aida02 ~ 154

aida05 ~ 27

aida03 ~ 181

aida04 ~ 193

Moved power cable for 19 inch rack to CRYRING 'clean' power source

aida03 ~ 211

aida04 ~ 212

aida01 ~ 317

aida02 ~ 215

aida05 ~ 29

Reconnected CRYRING ground 

aida01 ~ 301

aida02 ~ 203

aida05 ~ 28

aida03 ~ 181

aida04 ~ 191

Current status 17:00:
All CARME vacuum equipment is off (turbo, SIP power modules, NEG power modules, gauges) - roughing pump is on but is valved off from the turbo which is valved off from chamber

Motors and motors pi are off. Detectors are fully out and the pin has been reinserted. 

Bias ON 100V, 8.5 uA. Pulser ON, all equipment connected to 19 inch rack power ON.  Jumper is on link 1 for nn bias adaptor card only. 

Attachment 1: 19.PNG  380 kB  Uploaded Thu Aug 11 21:18:43 2022  | Hide | Hide all
19.PNG
Attachment 2: 18.PNG  377 kB  Uploaded Thu Aug 11 21:18:43 2022  | Hide | Hide all
18.PNG
Attachment 3: 17.PNG  259 kB  Uploaded Thu Aug 11 21:18:43 2022  | Hide | Hide all
17.PNG
Attachment 4: 16.PNG  261 kB  Uploaded Thu Aug 11 21:18:43 2022  | Hide | Hide all
16.PNG
Attachment 5: 15.PNG  263 kB  Uploaded Thu Aug 11 21:18:43 2022  | Hide | Hide all
15.PNG
Attachment 6: 14.PNG  200 kB  Uploaded Thu Aug 11 21:18:43 2022  | Hide | Hide all
14.PNG
Attachment 7: 13.PNG  365 kB  Uploaded Thu Aug 11 21:18:43 2022  | Hide | Hide all
13.PNG
Attachment 8: 12.PNG  371 kB  Uploaded Thu Aug 11 21:18:43 2022  | Hide | Hide all
12.PNG
Attachment 9: 11.PNG  260 kB  Uploaded Thu Aug 11 21:18:43 2022  | Hide | Hide all
11.PNG
Attachment 10: 10.PNG  259 kB  Uploaded Thu Aug 11 21:18:43 2022  | Hide | Hide all
10.PNG
Attachment 11: 9.PNG  264 kB  Uploaded Thu Aug 11 21:18:43 2022  | Hide | Hide all
9.PNG
Attachment 12: 8.PNG  290 kB  Uploaded Thu Aug 11 21:18:43 2022  | Hide | Hide all
8.PNG
Attachment 13: 7.PNG  216 kB  Uploaded Thu Aug 11 21:18:43 2022  | Hide | Hide all
7.PNG
Attachment 14: 6.PNG  204 kB  Uploaded Thu Aug 11 21:18:43 2022  | Hide | Hide all
6.PNG
ELOG V3.1.4-unknown